OpenStudy (anonymous):

I was wondering why a tractor has huge rear wheels but small front wheels ! Large wheels provide better grip (friction) with the ground. So, why are front wheels so small?

OpenStudy (fifciol):

because you need to turn left and right so it's easier to do that with small wheels(smaller inertia, smaller friction)

OpenStudy (theeric):

I agree with Fifciol. And I also think they're small so that they're economical due to lower cost. In some tractors, small wheels mean you can see more. (((Before you read the next paragraph, you should know that the great advantage to large wheels is the smaller pressure, \(P=\dfrac{F}{A}\), and so it will not exert as large pressures to the soil, which would cause the soil to move, leaving the tractor pressing into the ground. The soil might be loose, or even muddy.))) Large wheels are usually not as necessary, because I think they carry a lesser weight, and so a smaller contact area does not mean the pressure is greater than the pressure on the rear tires. Remember that pressure is defined as \(P=\dfrac{F}{A}\) and can be applied generally to this scenario. Let's look at the back tire.\[\frac{F_{w,\ rear}}{A_{rear}}=P_{rear}\]If you look at the changes from back tire to front tire - decreased weight and area - how will the pressure be diffent? If the weight and area decrease proportionally, pressure will stay the same. If they're decreased nearly proportionally, there will be a slight change in pressure.\[\frac{\left(F_{w,\ rear}\downarrow \right)}{\left(A_{rear}\downarrow\right)} = P_{rear}\updownarrow\] I would think that you'd need lower pressure on the back wheels, anyway, because that is where the soil is pushed back to get the tractor moving (Newton's Third Law). So the tires are able to be smaller because they don't have to be as large. And then, as Fifciol pointed out, that is good because of the lesser linear and rotational inertias.

OpenStudy (anonymous):

ok thanks